# Spinning black holes in astrophysical environments

### v1

@@ Abstract
Introduction. Vacuum black hole (BH) solutions are idealized models. In reality, astrophysical compact objects exist within complex matter-rich environments, including accretion disks and dark matter halos. Understanding how these environments influence a BH, affect observational signatures, and offer insights into the nature of the surrounding matter is a growing field of study. To investigate how the astrophysical environment influences the BH geometry, one must solve the Einstein field equations with a stress-energy tensor that captures the environmental effects. In recent years, the Einstein cluster [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] has been used as a model for dark matter halos [12, 13, 14, 15, 16, 17], and to study their impact on BH geometries [18, 19, 20, 21, 16, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17].
@@
where $\varepsilon$, $p_{r}$, $p_{1}$ and $p_{2}$ are the energy density, radial pressure and transverse pressures in the comoving frame of the fluid, respectively. The vectors $u^{\mu}$, $k^{\mu}$ and $s^{\mu}$ are, respectively, the four-velocity of the fluid and spacelike vectors that define the directions of anisotropy such that $u^{\mu}u_{\mu}=-1$, $k^{\mu}k_{\mu}=1$, $s^{\mu}s_{\mu}=1$, $u^{\mu}k_{\mu}=0$, $u^{\mu}s_{\mu}=0$, and $k^{\mu}s_{\mu}=0$. The eigenvalues of the stress-energy tensor are $-\varepsilon$, $p_{r}$, $p_{1}$ and $p_{2}$.
To obtain rotating BH solutions, we consider a generic line-element that is stationary, axially-symmetric and circular. We use quasi-isotropic coordinates and the metric is described by four functions, $f$, $g$, $h$, $\omega$, of $r$ and $\theta$
@@
where $\kappa$ is the surface gravity on the event horizon, $A_{H}$ the area of the event horizon, $T$ the trace of the stress-energy tensor, and $J$ is the total angular momentum given by $J=J_{H}+\int d^{3}x\sqrt{-g}T^{t}_{\phantom{t}\varphi}$, where $J_{H}$ is the angular momentum of the BH, computed as a Komar integral. In line with the standard properties of the Einstein cluster, we take the dark matter halo to not have any contribution to the net angular momentum, such that all angular momentum is due to the BH’s rotation, $J=J_{H}$. This occurs when $T^{t}_{\phantom{t}\varphi}=0$, which is true when we choose the rotation law $\Omega=\omega$, i.e., the angular velocity of the dark matter halo is the frame-dragging angular velocity and therefore it has zero velocity with respect to the ZAMO. Note that the synchronization condition at the horizon is respected in this case. Under these conditions, Eq. (5) can be rewritten as $M=M_{\rm BH}+M_{\rm halo}$, where $M_{\rm BH}=\frac{\kappa}{4\pi}A_{H}+2\Omega_{H}J_{H}$, is the mass of the BH, and $M_{\rm halo}$ corresponds to the integral in Eq. (5). The total mass $M$ and angular momentum $J$ can also be computed from the asymptotic decay of the metric components $g_{tt}=-1+\frac{2M}{r}+\mathcal{O}(r^{-2})$, and $g_{t\varphi}=-\frac{2J}{r}\sin^{2}\theta+\mathcal{O}(r^{-2})$, while the surface gravity and the area of the event horizon for the metric (2) are computed as
@@
Solving the field equations. To solve the field equations and obtain stationary, axially-symmetric BH solutions, we follow the method of Ref. [51]. There are six non-trivial components of the Einstein field equations, namely the $tt$, $rr$, $\theta\theta$, $\varphi\varphi$, $t\varphi$ and $r\theta$ components. These are split into two groups. Defining $E_{\mu\nu}=G_{\mu\nu}-8\pi T_{\mu\nu}$, we consider the following combinations of field equations: $E_{\mu}^{\phantom{\mu}\mu}-2E_{r}^{\phantom{r}r}-2\omega E_{\varphi}^{\phantom{\varphi}t}=0$, $E_{\varphi}^{\phantom{\varphi}t}=0$, $E_{r}^{\phantom{r}r}+E_{\theta}^{\phantom{\theta}\theta}=0$, and $E_{\varphi}^{\phantom{\varphi}\varphi}-\omega E_{\varphi}^{\phantom{\varphi}t}-E_{r}^{\phantom{r}r}-E_{\theta}^{\phantom{\theta}\theta}=0$. These combinations ensure that each equation contains second-derivatives of only one of the four metric functions. The remaining linearly independent equations are taken to be $E_{r}^{\phantom{r}\theta}=0$, and $E_{r}^{\phantom{r}r}-E_{\theta}^{\phantom{\theta}\theta}=0$. However, as explained in Refs. [52, 53], these are directly related to the conservation of the stress-energy tensor, $\nabla_{\mu}T^{\mu\nu}=0$, via the Bianchi identities, and therefore can be substituted by the conditions  [^2] In the absence of rotation, the system becomes spherically symmetric, and the condition $\nabla_{\mu}T^{\mu\theta}=0$, automatically implies $p_{1}=p_{2}$. $\nabla_{\mu}T^{\mu r}=0$, and $\nabla_{\mu}T^{\mu\theta}=0$. The system contains eight undetermined functions of $r$ and $\theta$, namely $f$, $g$, $h$, $\omega$, $\varepsilon$, $p_{r}$, $p_{1}$ and $p_{2}$. The four combinations of Einstein equations, together with the conservation of the stress-energy tensor, determine six of these functions, namely the four metric functions and the two tangential pressures $p_{1}$ and $p_{2}$. To close the system we choose a profile for the energy density $\varepsilon$ of the dark matter halo, and provide an equation of state relating $p_{r}$ and $\varepsilon$. For the equation of state we choose $p_{r}=0$, in accordance with the Einstein cluster. Motivated by the Hernquist profile [55], and by Ref. [18], we use the following profile for the energy density
@@
where $a_{0}$ is a typical length scale associated with the halo, and $\mathcal{M}$ represents the mass of the halo $M_{\rm halo}$ to leading order in $\mathcal{M}/a_{0}$ in the static limit. The function $b\equiv b(r)$ is a complicated function of $r$ presented in the Supplemental Material, and as explained there, it guarantees that the parameters $\mathcal{M}$ and $a_{0}$ have suitable physical meaning. Importantly, $b>0$, and thus $\varepsilon\geq 0$ for physically relevant choices of the parameters $\mathcal{M}$ and $a_{0}$. In the large $r$ limit, this profile for $\varepsilon$ agrees with the Hernquist profile.
The functions obey the following boundary conditions. At the horizon ($r=r_{H}$) we have $\partial_{r}f=\partial_{r}g=\partial_{r}h=\partial_{r}p_{1}=\partial_{r}p_{2}=0,\quad\omega=\Omega_{H}$, while asymptotically they obey $f=g=h=1,\quad\omega=p_{1}=p_{2}=0$. Regularity, axial symmetry and parity considerations imply $\partial_{\theta}f=\partial_{\theta}g=\partial_{\theta}h=\partial_{\theta}\omega=\partial_{\theta}p_{1}=\partial_{\theta}p_{2}=0$, at $\theta=0$ and $\theta=\pi/2$  [^3] Incidentally, the field equations impose that $p_{1}=p_{2}$, at $\theta=0$. We do not impose this relation in the numerical method, but use it as a test to the code.. The input parameters are $r_{H}$, $\Omega_{H}$, and the parameters describing the dark matter halo, $a_{0}$ and $\mathcal{M}$.
@@
Results. The parameter space is vast: fixing the BH mass, it is determined by the BH spin, the mass of the halo, and the length scale $a_{0}$. Increasing $a_{0}$ while keeping the other parameters constant significantly complicates the numerical problem. For sufficiently large $a_{0}$ (low halo compactness $M_{\rm halo}/a_{0}$), the numerical method requires very high resolution due to steep gradients in the metric functions – caused by the dark matter halo – emerging within a narrow region of the radial coordinate. For these reasons, we have focused on the case $M_{\rm halo}\approx 10M_{\rm BH}$ and $a_{0}\approx 100M_{\rm BH}$, ensuring a clear hierarchy between $M_{\rm BH}$, $M_{\rm halo}$, and $a_{0}$, while keeping the computational cost manageable for a systematic study. This setup is sufficient for our primary objective: to investigate the impact of spin on BHs in astrophysical environments. We have also examined solutions with $M_{\rm halo}\approx 10M_{\rm BH}$ and $a_{0}\approx 1000M_{\rm BH}$, finding that the qualitative conclusions remain consistent.
@@
First, we examined whether rotation induces violations of the standard energy conditions by studying the eigenvalues of the stress-energy tensor. The energy conditions are as follows. Null energy condition: $\varepsilon+p_{i}\geq 0$; Weak energy condition: $\varepsilon\geq 0$ and $\varepsilon+p_{i}\geq 0$; Strong energy condition: $\varepsilon+p_{i}\geq 0$ and $\varepsilon+\sum_{i}p_{i}\geq 0$; Dominant energy condition: $\varepsilon\geq|p_{i}|$. Here, $i$ runs over $\{r,1,2\}$. In the static case originally studied in Ref. [18], it was found that only the dominant energy condition is violated, in a small region close to the horizon. However, simple modifications of the original profile, e.g. by using a cutoff radius in the profile for the energy density, eliminate this issue [18, 20, 23, 22]. Since, for simplicity, we do not impose such a cutoff radius in our profile for the energy density, Eq. (7), we expect similarly to Ref. [18], violations of the dominant energy condition in a small region close to the horizon. In Fig. 1, we show the equatorial plane profiles for the eigenvalues of the stress-energy tensor for a solution with $J/M_{\rm BH}^{2}\approx 0.8$. All eigenvalues remain non-negative throughout the domain, indicating that the null, weak, and strong energy conditions are satisfied. As anticipated, the dominant energy condition is violated near the horizon. Similar profiles are observed for other values of $\theta$, and the same conclusions apply to solutions with different spins. Thus, once a suitable modified profile for the energy density is used, we expect that solutions in this framework to obey all energy conditions.
@@
Acknowledgments. We thank Kyriakos Destounis, Andrea Maselli, José Natário, and Prashant Kocherlakota for useful comments on a previous version of the manuscript. P.F. would like to thank CENTRA, Instituto Superior Técnico, for their kind hospitality during the initial stages of this work. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. We acknowledge support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. V.C. is a Villum Investigator and a DNRF Chair. V.C. acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101007855 and No 101131233. P.F. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).

### v2

@@ Abstract
Introduction. Vacuum black hole (BH) solutions are idealized models. In reality, astrophysical compact objects exist within complex matter-rich environments, including accretion disks and dark matter halos [1, 2]. Understanding how these environments influence a BH, affect observational signatures, and offer insights into the nature of the surrounding matter is a growing field of study. To investigate how the astrophysical environment influences the BH geometry, one must solve the Einstein field equations with a stress-energy tensor that captures the environmental effects. In recent years, the Einstein cluster [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] has been used as a model for dark matter halos [14, 15, 16, 17, 18, 19], and to study their impact on BH geometries [20, 21, 22, 23, 18, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 19, 37]. The Einstein cluster models a large collection of non-interacting particles moving along circular geodesics in all directions under their collective gravitational field. The emergent system is described by an anisotropic fluid with zero radial pressure.
@@
where $\varepsilon$, $p_{r}$, $p_{1}$ and $p_{2}$ are the energy density, radial pressure and transverse pressures in the comoving frame of the fluid, respectively. The vectors $u^{\mu}$, $k^{\mu}$ and $s^{\mu}$ are, respectively, the four-velocity of the fluid and spacelike vectors that define the directions of anisotropy such that $u^{\mu}u_{\mu}=-1$, $k^{\mu}k_{\mu}=1$, $s^{\mu}s_{\mu}=1$, $u^{\mu}k_{\mu}=0$, $u^{\mu}s_{\mu}=0$, and $k^{\mu}s_{\mu}=0$. The eigenvalues of the stress-energy tensor are $\varepsilon$, $p_{r}$, $p_{1}$ and $p_{2}$. This ansatz guarantees that there are no energy and momentum fluxes in the rest frame of the fluid. Such a fluid, modeling a dark matter halo in our case, is expected to generically be present near black holes as a result of accretion. As noted in Refs. [1, 2], the density profile produced by the (adiabatic) accretion-driven growth of a black hole typically develops an overdense cusp with a sharp cutoff in the vicinity of the horizon.
To obtain rotating BH solutions, we consider a line-element that is stationary, axially-symmetric and circular, belonging to the Weyl-Lewis-Papapetrou class. We use quasi-isotropic coordinates and the metric is described by four functions, $f$, $g$, $h$, $\omega$, of $r$ and $\theta$
@@
where $\kappa$ is the surface gravity on the event horizon, $A_{H}$ the area of the event horizon, $T$ the trace of the stress-energy tensor, and $J$ is the total angular momentum given by $J=J_{H}+\int d^{3}x\sqrt{-g}T^{t}_{\phantom{t}\varphi}$, where $J_{H}$ is the angular momentum of the BH, computed as a Komar integral. We adopt a model for a dark matter halo similar to the Einstein cluster, where the dark matter halo has zero angular momentum. Consequently, the total angular momentum of the system is $J=J_{H}$. This occurs when $T^{t}_{\phantom{t}\varphi}=0$, which is true when we choose the rotation law $\Omega=\omega$, i.e., the angular velocity of the dark matter halo is the frame-dragging angular velocity and therefore it has zero velocity with respect to the ZAMO. Note that the synchronization condition at the horizon is respected in this case. Under these conditions, Eq. (5) can be rewritten as $M=M_{\rm BH}+M_{\rm halo}$, where $M_{\rm BH}=\frac{\kappa}{4\pi}A_{H}+2\Omega_{H}J_{H}$, is the mass of the BH, and $M_{\rm halo}$ corresponds to the integral in Eq. (5). The total mass $M$ and angular momentum $J$ can also be computed from the asymptotic decay of the metric components $g_{tt}=-1+\frac{2M}{r}+\mathcal{O}(r^{-2})$, and $g_{t\varphi}=-\frac{2J}{r}\sin^{2}\theta+\mathcal{O}(r^{-2})$, while the surface gravity and the area of the event horizon for the metric (2) are computed as
@@
Solving the field equations. To solve the field equations and obtain stationary, axially-symmetric BH solutions, we follow the method of Ref. [54]. There are six non-trivial components of the Einstein field equations, namely the $tt$, rrrr, $\theta\theta$, $\varphi\varphi$, $t\varphi$ and $r\theta$ components. These are split into two groups. Defining $E_{\mu\nu}=G_{\mu\nu}-8\pi T_{\mu\nu}$, we consider the following combinations of field equations: $E_{\mu}^{\phantom{\mu}\mu}-2E_{r}^{\phantom{r}r}-2\omega E_{\varphi}^{\phantom{\varphi}t}=0$, $E_{\varphi}^{\phantom{\varphi}t}=0$, $E_{r}^{\phantom{r}r}+E_{\theta}^{\phantom{\theta}\theta}=0$, and $E_{\varphi}^{\phantom{\varphi}\varphi}-\omega E_{\varphi}^{\phantom{\varphi}t}-E_{r}^{\phantom{r}r}-E_{\theta}^{\phantom{\theta}\theta}=0$. These combinations ensure that each equation contains second-derivatives of only one of the four metric functions. The remaining linearly independent equations are taken to be $E_{r}^{\phantom{r}\theta}=0$, and $E_{r}^{\phantom{r}r}-E_{\theta}^{\phantom{\theta}\theta}=0$. However, as explained in Refs. [55, 56], these are directly related to the conservation of the stress-energy tensor, $\nabla_{\mu}T^{\mu\nu}=0$, via the Bianchi identities, and therefore can be substituted by the conditions $\nabla_{\mu}T^{\mu r}=0$, and $\nabla_{\mu}T^{\mu\theta}=0$. The system contains eight undetermined functions of $r$ and $\theta$, namely $f$, $g$, $h$, $\omega$, $\varepsilon$, $p_{r}$, $p_{1}$ and $p_{2}$. The four combinations of Einstein equations, together with the conservation of the stress-energy tensor, determine six of these functions, that we choose to be the four metric functions and the two tangential pressures $p_{1}$ and $p_{2}$  [^2] In the spherically symmetric limit, the field equations impose $p_{1}=p_{2}$. To close the system we choose a profile for the energy density $\varepsilon$ of the dark matter halo, and provide an equation of state relating $p_{r}$ and $\varepsilon$. For the equation of state we choose $p_{r}=0$, in accordance with the Einstein cluster. Motivated by the Hernquist profile [58], and by Ref. [20], we use the following profile for the energy density
@@
where $a_{0}$ is a typical length scale associated with the halo, and $\mathcal{M}$ represents the mass of the halo $M_{\rm halo}$ to leading order in $\mathcal{M}/a_{0}$ in the static limit. The function $b\equiv b(r)$ is a complicated function of $r$ presented in the Supplemental Material, and as explained there, it guarantees that the parameters $\mathcal{M}$ and $a_{0}$ have suitable physical meaning. Importantly, $b>0$, and thus $\varepsilon\geq 0$ for physically relevant choices of the parameters $\mathcal{M}$ and $a_{0}$. For large $r$, as well as in the absence of a black hole, this profile for the energy density agrees with the Hernquist profile. This energy density profile therefore models a dark matter halo co-existing with a black hole.
The functions obey the following boundary conditions. At the horizon ($r=r_{H}$), by solving the field equations order-by-order in a near-horizon expansion, we find $\partial_{r}f=\partial_{r}g=\partial_{r}h=\partial_{r}p_{1}=\partial_{r}p_{2}=0,\quad\omega=\Omega_{H}$, while asymptotically they obey $f=g=h=1,\quad\omega=p_{1}=p_{2}=0$. Regularity, axial symmetry and parity considerations imply $\partial_{\theta}f=\partial_{\theta}g=\partial_{\theta}h=\partial_{\theta}\omega=\partial_{\theta}p_{1}=\partial_{\theta}p_{2}=0$, at $\theta=0$ and $\theta=\pi/2$  [^3] Incidentally, the field equations impose that $p_{1}=p_{2}$, at $\theta=0$. We do not impose this relation in the numerical method, but use it as a test to the code.. The input parameters are $r_{H}$, $\Omega_{H}$, and the parameters describing the dark matter halo, $a_{0}$ and $\mathcal{M}$.
@@
Results. The parameter space is vast: fixing the BH mass, it is determined by the BH spin, the mass of the halo, and the length scale $a_{0}$. Increasing $a_{0}$ while keeping the other parameters constant significantly complicates the numerical problem. For sufficiently large $a_{0}$ (low halo compactness $M_{\rm halo}/a_{0}$), the numerical method requires very high resolution due to steep gradients in the metric functions emerging within a narrow region of the numerical grid. This occurs because around the characteristic length scale, $r\sim a_{0}$ (large values of $r$), the energy density profile transitions between different regimes: $\varepsilon\sim 1/r$ when $r\ll a_{0}$, while $\varepsilon\sim 1/r^{4}$ when $r\gg a_{0}$, as is characteristic of the Hernquist profile. We discuss this issue in more detail in the Supplemental Material. For these reasons, we have focused on the case $M_{\rm halo}\approx 10M_{\rm BH}$ and $a_{0}\approx 100M_{\rm BH}$, ensuring a clear hierarchy between $M_{\rm BH}$, $M_{\rm halo}$, and $a_{0}$, while keeping the computational cost manageable for a systematic study. This setup is sufficient for our primary objective: to investigate the impact of spin on BHs in astrophysical environments. We have also examined solutions with $M_{\rm halo}\approx 10M_{\rm BH}$ and $a_{0}\approx 1000M_{\rm BH}$, finding that the qualitative conclusions remain consistent.
@@
First, we examined whether rotation induces violations of the standard energy conditions by studying the eigenvalues of the stress-energy tensor. The energy conditions are as follows. Null energy condition: $\varepsilon+p_{i}\geq 0$; Weak energy condition: $\varepsilon\geq 0$ and $\varepsilon+p_{i}\geq 0$; Strong energy condition: $\varepsilon+p_{i}\geq 0$ and $\varepsilon+\sum_{i}p_{i}\geq 0$; Dominant energy condition: $\varepsilon\geq|p_{i}|$. Here, $i$ runs over $\{r,1,2\}$. In the static case originally studied in Ref. [20], it was found that only the dominant energy condition is violated, in a small, almost empty, region close to the horizon. However, simple modifications of the original profile, e.g. by using a cutoff radius in the profile for the energy density, eliminate this issue [20, 22, 25, 24]. Since, for simplicity, we do not impose such a cutoff radius in our profile for the energy density, Eq. (7), we expect similarly to Ref. [20], violations of the dominant energy condition in a small region close to the horizon. However, these modifications of the energy-density profile are not expected to significantly affect the physical solution, since they occur in an almost empty region with negligible energy density and therefore have no influence on the spacetime dynamics. In Fig. 1, we show the equatorial plane profiles for the eigenvalues of the stress-energy tensor for a solution with $J/M_{\rm BH}^{2}\approx 0.8$. All eigenvalues remain non-negative throughout the domain, indicating that the null, weak, and strong energy conditions are satisfied. As anticipated, the dominant energy condition is violated near the horizon. Similar profiles are observed for other values of $\theta$, and the same conclusions apply to solutions with different spins. Thus, once a suitable modified profile for the energy density is used, we expect that solutions in this framework to obey all energy conditions.
@@
Acknowledgments. We thank Kyriakos Destounis, Andrea Maselli, José Natário, and Prashant Kocherlakota for useful comments on a previous version of the manuscript. P.F. would like to thank CENTRA, Instituto Superior Técnico, for their kind hospitality during the initial stages of this work. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. DNRF184. We acknowledge support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. V.C. is a Villum Investigator and a DNRF Chair. V.C. acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101007855 and No 101131233. P.F. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).
@@
Numerical issues for large characteristic length scale $a_{0}$
The scale $a_{0}$ marks the characteristic scale at which the energy density of the halo transitions from $\varepsilon\sim 1/r$ when $r\ll a_{0}$, to $\varepsilon\sim 1/r^{4}$ when $r\gg a_{0}$. This leads to steep gradients in the metric functions, which are difficult to capture with the numerical method.
As an example, we consider the static limit, where we know the metric function $b(r)^{4}=fN_{+}^{4}/g$ analytically, as obtained in the sections above. In Fig. 4 we plot $b^{\prime}(x)$ as a function of $x=1-2r_{H}/r$, which is the radial coordinate employed numerically. We observe that when $a_{0}$ is large, there are steep gradients in the metric functions, which appear in a small range of $x$, that a numerical method struggles to capture accurately. For this reason, for large values of $a_{0}$, a (very) large resolution in $x$ must be used.
Figure 4: $b^{\prime}(x)$ as a function of $x=1-2r_{H}/r$, for $M_{\rm halo}=10r_{H}$, and two distinct values of $a_{0}$. We observe steep gradients when $a_{0}$ is large. For comparison, in the vacuum GR case, b(x)=(3x)/2b(x)=(3-x)/2, such that $b^{\prime}(x)=-1/2$.